OpenStudy (anonymous):

What shortcut (formula) is applicable for simplifying the following expression? x^4 times x^7 Divide the exponents when you have the same base. Multiply the exponents when you have the same base. Add the exponents when you have the same base. Subtract the exponents when you have the same base.

OpenStudy (lucaz):

for example, (x^3)*(x^2) = (x*x*x)*(x*x) = x^5 = x^(3+2)

OpenStudy (lucaz):

sorry, it's x^(4+7) = x^11 not 13

OpenStudy (anonymous):

so the answer is B ?